# FormulaReasoning: A Dataset for Formula-Based Numerical Reasoning

### v5

@@ Abstract
The application of formulas (e.g., physics formulas) is a fundamental ability of humans when solving numerical reasoning problems. Existing numerical reasoning datasets seldom explicitly indicate the formulas employed in reasoning, as their questions rely on implicit commonsense mathematical knowledge. In contrast, in this paper, we introduce FormulaReasoning, a new dataset specifically designed for formula-based numerical reasoning. Each of the 4,751 questions in our dataset requires numerical calculation with external physics formulas, making it a more challenging benchmark for evaluating large language models (LLMs). We offer normalized fine-grained annotations for the questions, available in English and Chinese, including formula structures, parameter names, symbols, numerical values, and units, derived from extensive manual effort with LLM assistance for guaranteed quality. We also provide a consolidated formula database to serve as an external knowledge base accompanying the dataset. We employ FormulaReasoning to evaluate LLMs with 7B to over 100B parameters, and explore retrieval-augmented generation with the formula database. Our evaluation also covers supervised methods that break down the reasoning process into formula generation, parameter extraction, and numerical calculation, as well as direct preference optimization methods based on derived preference data.
@@ 1 Introduction
Numerical reasoning constitutes one of the significant forms within natural language reasoning (Frieder et al., 2023). The study of numerical reasoning has seen substantial progress in recent years, driven largely by the development of LLMs (OpenAI, 2023; Touvron et al., 2023; Yang et al., 2024a) and specialized datasets (Wang et al., 2017; Dua et al., 2019; Amini et al., 2019; Cobbe et al., 2021a). Current datasets for numerical reasoning typically include simple commonsense numerical questions that under-reflect the complexity of real-world problems. These datasets either do not provide process supervision information, or the provided reasoning steps are essentially incomplete, as they often rely on implicit commonsense knowledge but not explicit knowledge guiding the reasoning process. This issue becomes particularly evident when LLMs experience hallucination (Frieder et al., 2023; Bang et al., 2023), especially in the absence of clear knowledge guidance. Consequently, one might ask “What explicit knowledge could be used to guide a numerical reasoning process?”. Formulas exactly represent such knowledge that has been largely overlooked in previous research but is frequently utilized in real-life applications.
@@
Take a question from the GSM8K dataset (Cobbe et al., 2021a) as an example: “A robe takes 2 bolts of blue fiber and half that much white fiber. How many bolts in total does it take?” This question implicitly uses simple commonsense mathematical knowledge (e.g., half means dividing by 2) to solve without requiring complex domain-specific formular knowledge for numerical reasoning. Recently, Liu et al. Liu et al. (2023) constructed two formula-based datasets, Math23K-F and MAWPS-F. However, only 33% and 38% of the questions in these datasets, respectively, require the use of formulas and still mainly consist of simple commonsense formulas (e.g., total_amount = unit_amount $\times$ total_number). Currently, there is a lack of datasets where questions require domain-specific formulas to guide complex numerical reasoning, such as the physics formula used to calculate the heat absorption of an object.
@@
To fill the gap, we constructed FormulaReasoning, a new dataset for numerical reasoning that requires the use of domain-specific formulas (i.e., physics). We annotated formulas for each question as illustrated in Figure 1. This formula-based feature makes FormulaReasoning a challenging dataset to evaluate systems with real-world complex numerical reasoning problems. In Table 1, we compare FormulaReasoning with two existing formula-based datasets and the well-known formula-free datasets GSM8K and MATH (Hendrycks et al., 2021). Compared to Math23K-F and MAWPS-F, FormulaReasoning contains more diverse formulas (272 vs. 18–51), and a larger proportion of questions requiring formulas (100% vs. 33–38%). Moreover, the higher average number of reasoning steps (2.39 vs. 1.01–1.16) implies that FormulaReasoning is more challenging and better suited to evaluate formula-based multi-step numerical reasoning.
@@
Dataset Math23K-F MAWPS-F GSM8K MATH FormulaReasoning # questions 23,162 2,373 8,792 12,500 4,751 # formulas (and variants) 51 (131) 18 (46) 0 (0) 0 (0) 272 (824) # questions requiring formulas (proportion) 7,750 (33%) 911 (38%) N/A N/A 4,751 (100%) Avg. # reasoning steps 1.16 1.01 3.59 Not Provided 2.39
@@
We used FormulaReasoning to evaluate LLMs with 7B to $>$100B parameters, as well as small models fine-tuned with proposed Chain-of-Thought (CoT) supervision and data augmentation. We also trained an encoder for formula retrieval and experimented with retrieval-augmented generation (RAG). In addition, we derived preference data and evaluated direct preference optimization (DPO) methods. Our empirical findings show that the best existing models only achieve an accuracy of 78%, indicating that there is significant room for research in formula-based numerical reasoning.
@@
  • We perform an evaluation of LLMs of various sizes, RAG, fine-tuned small models, and DPO methods based on derived preference data. Our experimental results establish a solid baseline for future research and also indicate that the task remains an unsolved challenge.
@@ 2.1 Numerical Reasoning Datasets
The study of numerical reasoning in natural language has existed for years. Numerous datasets, such as DROP (Dua et al., 2019), GSM8K (Cobbe et al., 2021b), TSQA (Li et al., 2021), and MATH (Hendrycks et al., 2021), have addressed natural language numerical reasoning. Another line of research that focuses on numerical reasoning in natural language is math word problem (MWP). MWP tasks typically provide a short passage (i.e., a question) and require the generation of an arithmetic expression that can compute an answer. Representative datasets include MAWPS (Koncel-Kedziorski et al., 2016), Math23K (Wang et al., 2017), MathQA (Amini et al., 2019), etc. Several works focus on numerical reasoning in specialized domains. Examples include GeoSQA (Huang et al., 2019), which focuses on the geography domain, STEM (Drori et al., 2023) and ScienceQA (Lu et al., 2022) which cover multiple disciplines in science and technology. Distinguished from these datasets, the numerical reasoning questions in our FormulaReasoning have explicitly labeled formulas.
The recently introduced datasets Math23K-F and MAWPS-F (Liu et al., 2023) require formulas for only 33% and 38% of the questions, respectively, and their formulas are all simple commonsense formulas (e.g., total_cost = unit_cost $\times$ total_number). By contrast, our FormulaReasoning adapts questions from physics exams, with every question accompanied by a fine-grained annotation of domain-specific formulas. In addition, we provide a consolidated formula database that can serve as an external knowledge base to assess RAG solutions.
@@ 2.2 Numerical Reasoning Methods
Our experiments implemented and evaluated representative existing methods as baselines for FormulaReasoning, including zero-shot CoT prompting for LLMs ranging from 7B to more than 100B parameters. We trained a specialized formula retriever to be used in RAG. We divided the reasoning process into formula generation, parameter extraction, and numerical calculation, and used data augmentation to enhance fine-tuned small models. We derived preference data and performed DPO.
@@ 3 Dataset Construction
We collected questions from Chinese junior high school physics exams. We invited five graduate students to act as annotators, all of whom hold a bachelor’s degree in science and engineering. For all questions, we normalized their explanations. Each question was annotated with normalized reasoning steps in natural language and a tabular representation of the steps using formulas, including the numerical values and units for all the parameters in the formulas. This process involved a combination of manual annotation and the assistance of an LLM [^1] During dataset construction, we consistently accessed Qwen-max as our LLM via its API https://help.aliyun.com/zh/dashscope/developer-reference/quick-start. to improve the efficiency of annotation. We compiled all the formulas and merged those that express the same meaning to create a consolidated formula database. Finally, we altered the questions to avoid label leakage and translated the dataset into English. In the following, we elaborate this process to construct FormulaReasoning.
@@ 3.2 Formula Normalization
The reasoning steps in the raw explanation text lacked a normalized format and were expressed casually. Some formulas mixed parameter names (e.g., “mass of water”) with symbols (e.g., “$m_{water}$”), while others simply provided calculations over numerical values without parameter names or symbols. To ensure that all explanations adopted a common form of formulas, we performed normalization, as illustrated in Table 2. We aim to identify and normalize the formulas in the original explanations and then verify and unify their formats. Manually carrying out such tasks would require a significant effort. However, since the process is not open-ended, but rather structured and verifiable, we could automatically, e.g. using an LLM, extract formulas from a normalized explanation, calculate each step, and compare the result with the given answer to ensure the accuracy of our normalization.
Specifically, to improve the efficiency of annotation, we adopted a coarse-to-fine annotation process with the help of an LLM. We first prompted the LLM to revise the reasoning steps (in particular, the formulas) in the explanation text in a normalized form. Then, we prompted the LLM to correct minor errors within the normalized explanations, including formatting issues in formula annotations and inaccuracies in the parameters used during calculation. In the following, we elaborate this two-stage process. More details, including the prompts, are provided in Appendix A.1.1.
@@
We introduced each question with its original explanation and answer to guide the LLM through a few-shot prompt to normalize the explanation. We observed that the quality of the normalized explanations was generally satisfactory. To assess the correctness of a normalized explanation, we extracted formulas from the explanation and calculated an answer using the Numbat calculator. [^2] https://numbat.dev. Numbat is designed for scientific computation with support for physical units. We also required the LLM to present the symbol, numerical value, and unit of each parameter in formulas in the form of a table, as illustrated in Figure 1.
@@
We checked the correctness of the formula format in the explanations by rules, including whether there were omissions in parameter names, symbols, or units, and these minor issues were correctable. If we detected that the LLM generated an ill-formatted normalized explanation, we used a few-shot prompt to identify and correct these specific errors. We observed that the questions that remained incorrect despite multiple attempts by the LLM were of inherently poor quality, e.g., missing important reasoning steps, as illustrated in Appendix A.1.2. These questions were removed from our dataset. After that, our dataset contains 5,420 questions remaining.
@@ 3.3 Formula Database Construction
We divided the construction process of a consolidated formula database into three steps: 1) Merge formulas through symbolic rules. 2) Merge formulas through a semantics-based method. 3) Review manually and correct errors. In Table 3, we present the initial number of formulas and the remaining number of formulas after each step.
@@
This was achieved by comparing formula structures and symbols. Consider the following as an example of judging whether two formulas have the same structure. The three formulas “$f_{1}:~{}a_{1}=(b_{1}+c_{1})/d_{1}$”, “$f_{2}:a_{2}=(b_{2}+c_{2})/d_{2}$”, and “$f_{3}:b_{1}=a_{1}*d_{1}-c_{1}$” have the same structure because $f_{2}$ can be derived from $f_{1}$ by renaming parameters, and $f_{3}$ can be obtained from $f_{1}$ by transformation. Moreover, in physics, physical quantities are conventionally represented by specific symbols. For example, the mass of an object is often denoted by “$m$” and the density of an object is frequently represented by “$\rho$”. Subscripts are then used to specify which specific object a physical quantity refers to, such as “$\rho_{water}$” for the density of water. Therefore, for each pair of formulas, we first computed all the transformations of each formula to obtain a set of all its variants. Then, we compared the formula structures in the two sets to determine if the two formulas shared a structure. If so, we checked whether their symbols, with subscripts removed, were identical. If so, we considered these two formulas to be mergeable. When merging, we retained the parameter with the shorter name of the two. After merging based on such symbolic rules, we reduced the number of formulas in the formula database from 12,906 to 1,163.
@@ 3.4 Question Alteration
We replaced a random parameter in the original question with a placeholder “[X]”, and added the original answer to the question. The new question then asked for the value of the masked parameter “[X]”. Due to the various expressions of parameters, some could not be easily found and replaced in the question by string matching, e.g., “$4.2\times 10^{6}J/h$” and “$1\times 10^{-3}dm^{3}$” could be written as “$4.2$×$10^{6}$ J per hour” and “1L”, respectively. We had to exclude such questions.
@@
To verify that the modified questions were solvable, we used SymPy to solve the equations. During the verification process, some equations could not be solved (e.g., SymPy cannot handle “%” and “1000kg/t”), so we removed these questions. The final number of the remaining questions is 4,751.
@@ 3.5 English Version of FormulaReasoning
We employed the LLM to translate Chinese questions, explanations, and formula database into English. The translation prompts are provided in Appendix A.1.4. To evaluate translation quality, we manually translated a sample of 50 questions and calculated the perplexity of the LLM-translated version and the manually translated version using Qwen models for comparison. As shown in Table 4, the perplexity of LLM translations was lower than that of human translations. This was because for the altered questions, humans tended to translate the content in a direct way, while the LLM expressed it more fluently. It suggested that the English version translated by the LLM exhibited high quality.
Table 4: Perplexity of LLM and human translations.
Model LLM Human
Qwen2.5-14B 6.55 7.32
Qwen2.5-7B 7.52 8.04
We compared the Chinese and translated English versions of the 50 sampled questions, categorizing translation quality into two types: 1) All information was accurately translated or there were insignificant differences in expression. 2) There were translation errors that would affect question answering. We found that only 6% of the questions fell into the second category where the LLM had a small chance of translating “[X]” in the alteration of the question to other vague statements without “[X]”.

### v6

@@ Abstract
The application of formulas (e.g., physics formulas) is a fundamental human ability in solving numerical reasoning problems. Existing numerical reasoning datasets rarely explicitly state the formulas employed, as their questions often rely on implicit commonsense mathematical knowledge. To address this gap, we introduce FormulaReasoning, a new dataset specifically designed for formula-based numerical reasoning. It consists of 5,324 questions that require numerical calculations grounded in external physics formulas. We provide normalized, fine-grained annotations in both English and Chinese, including formula structures, parameter names, symbols, numerical values, and units—curated through extensive manual effort with LLM-assisted validation to ensure high quality. Additionally, we offer a consolidated formula database to serve as an external knowledge source. We analyze various reasoning approaches on FormulaReasoning, with emphasis on comparative evaluation of different architectural and methodological frameworks. Our assessment includes retrieval-augmented methods, approaches that decompose reasoning into formula generation, parameter extraction, and numerical calculation, as well as optimization techniques using preference data. We identify key challenges in formula-based numerical reasoning that require further investigation across different reasoning paradigms, highlighting opportunities for methodological advancement.
FormulaReasoning: A Dataset for Formula-Based Numerical Reasoning
Xiao Li* and Bolin Zhuthanks: These authors contributed equally to this work. and Kaiwen Shi Sichen Liu and Yin Zhu and Yiwei Liu and Gong Chengthanks: Corresponding author State Key Laboratory for Novel Software Technology, Nanjing University, Nanjing, China {xiaoli.nju,bolinzhu}@smail.nju.edu.cn {kaiwenshi,sichenliu,yinzhu,ywliu}@smail.nju.edu.cn gcheng@nju.edu.cn
@@ 1 Introduction
Numerical reasoning constitutes one of the significant forms within natural language reasoning (Frieder et al., 2023). The study of numerical reasoning has seen substantial progress in recent years, driven largely by the development of LLMs (OpenAI, 2023; Touvron et al., 2023; Yang et al., 2024) and specialized datasets (Wang et al., 2017; Dua et al., 2019; Amini et al., 2019; Cobbe et al., 2021a). Current datasets for numerical reasoning typically include simple commonsense numerical questions that under-reflect the complexity of real-world problems. These datasets either do not provide process supervision information, or the provided reasoning steps are essentially incomplete, as they often rely on implicit commonsense knowledge but not explicit knowledge guiding the reasoning process. This issue becomes particularly evident when LLMs experience hallucination (Frieder et al., 2023; Bang et al., 2023), especially in the absence of clear knowledge guidance. Consequently, one might ask “What explicit knowledge could be used to guide a numerical reasoning process?”. Formulas exactly represent such knowledge that has been largely overlooked in previous research but is frequently utilized in real-life applications.
@@
Take a question from the GSM8K dataset (Cobbe et al., 2021a) as an example: “A robe takes 2 bolts of blue fiber and half that much white fiber. How many bolts in total does it take?” This question implicitly uses simple commonsense mathematical knowledge (e.g., half means dividing by 2) to solve without requiring complex domain-specific formular knowledge for numerical reasoning. Recently, Liu et al. Liu et al. (2023) constructed two formula-based datasets, Math23K-F and MAWPS-F. Although 33% and 38% of the questions in these datasets, respectively, require the use of formulas, they still mainly consist of simple commonsense formulas (e.g., total_amount = unit_amount $\times$ total_number). Currently, there is a lack of datasets where questions require domain-specific formulas to guide complex numerical reasoning, such as the physics formula used to calculate the heat absorption of an object.
@@
To fill the gap, we constructed FormulaReasoning, a new dataset for numerical reasoning that requires the use of domain-specific formulas (i.e., physics). We annotated formulas for each question as illustrated in Figure 1. This formula-based feature enables FormulaReasoning to effectively differentiate model capabilities across different scales, particularly highlighting the limitations of smaller language models in handling domain-specific formula applications. In Table 1, we compare FormulaReasoning with two existing formula-based datasets and the well-known formula-free datasets GSM8K and MATH (Hendrycks et al., 2021). Compared to Math23K-F and MAWPS-F, FormulaReasoning contains more diverse formulas (272 vs. 18–51). Moreover, the higher average number of reasoning steps (2.39 vs. 1.01–1.16) provides a more comprehensive evaluation framework for understanding how different model architectures handle multi-step formula-based reasoning.
@@
Dataset Math23K-F MAWPS-F GSM8K MATH FormulaReasoning # questions 23,162 2,373 8,792 12,500 5,324 # questions requiring formulas 7,750 911 N/A N/A 5,324 # formulas (and variants) 51 (131) 18 (46) 0 (0) 0 (0) 272 (845) Avg. # reasoning steps 1.16 1.01 3.59 Not Provided 2.36
@@
We conducted extensive experiments on FormulaReasoning to compare different reasoning paradigms, including standard LLM approaches with 8B to >>100B parameters, fine-tuned models with Chain-of-Thought (CoT) supervision, data augmentation methods, and retrieval-augmented generation (RAG) methods. In addition, we derived preference data and evaluated direct preference optimization (DPO) methods. Our analysis reveals significant differences in how various approaches handle formula-based reasoning.
@@
  • We perform an evaluation of LLMs of various sizes, RAG, fine-tuned small models, and DPO methods based on derived preference data. Our experimental results establish a solid baseline for future research and also indicate that the performance disparity across model scales presents new research opportunities for efficient formula-based reasoning.
@@ 2.1 Numerical Reasoning Datasets
The study of numerical reasoning in natural language has existed for years. Numerous datasets, such as GSM8K (Cobbe et al., 2021b), TSQA (Li et al., 2021), and MATH (Hendrycks et al., 2021), have addressed natural language numerical reasoning. Another line of research that focuses on numerical reasoning in natural language is math word problem (MWP). MWP tasks typically provide a short passage (i.e., a question) and require the generation of an arithmetic expression that can compute an answer. Representative datasets include MAWPS (Koncel-Kedziorski et al., 2016), Math23K (Wang et al., 2017), MathQA (Amini et al., 2019), etc. Several works focus on numerical reasoning in specialized domains. Examples include GeoSQA (Huang et al., 2019), which focuses on the geography domain, STEM (Drori et al., 2023) and ScienceQA (Lu et al., 2022) which cover multiple disciplines in science and technology. Distinguished from these datasets, the numerical reasoning questions in our FormulaReasoning have explicitly labeled formulas.
The recently introduced datasets Math23K-F and MAWPS-F (Liu et al., 2023) require formulas for 33% and 38% of the questions, respectively, but their formulas are all simple commonsense formulas (e.g., total_cost = unit_cost $\times$ total_number). By contrast, our FormulaReasoning adapts questions from physics exams, with every question accompanied by a fine-grained annotation of domain-specific formulas. In addition, we provide a consolidated formula database that can serve as an external knowledge base to assess RAG solutions.
@@ 2.2 Numerical Reasoning Methods
Our experiments implemented and evaluated representative existing methods as baselines for FormulaReasoning, including zero-shot CoT prompting for LLMs ranging from 8B to more than 100B parameters. We trained a specialized formula retriever to be used in RAG. We divided the reasoning process into formula generation, parameter extraction, and numerical calculation, and used data augmentation to enhance fine-tuned small models. We derived preference data and performed DPO.
@@ 3 Dataset Construction
We collected questions from Chinese junior high school physics exams. We invited five graduate students to act as annotators, all of whom hold a bachelor’s degree in science and engineering. For all questions, we normalized their explanations. Each question was annotated with normalized reasoning steps in natural language and a tabular representation of the steps using formulas, including the numerical values and units for all the parameters in the formulas. This process involved a combination of manual annotation and the assistance of an LLM [^1] During dataset construction, we consistently accessed Qwen-max as our LLM via its API. to improve the efficiency of annotation. We compiled all the formulas and merged those that express the same meaning to create a consolidated formula database. Finally, we altered the questions to avoid label leakage and translated the dataset into English. In the following, we elaborate this process to construct FormulaReasoning.
@@ 3.2 Formula Normalization
The reasoning steps in the raw explanation text lacked a normalized format and were expressed casually. Some formulas mixed parameter names (e.g., “mass of water”) with symbols (e.g., $m_{water}$”), while others simply provided calculations over numerical values without parameter names or symbols. To ensure that all explanations adopted a common form of formulas, we performed normalization, as illustrated in Table 2. We aimed to identify and normalize the formulas in the original explanations and then verify and unify their formats. Manually carrying out such tasks would require a significant effort. However, since the process is not open-ended, but rather structured and verifiable, we could automatically, e.g. using an LLM, extract formulas from a normalized explanation, symbolicly calculate each step, and compare the result with the given answer to ensure the accuracy of our normalization.
Specifically, to improve the efficiency of annotation, we adopted a coarse-to-fine annotation process with the help of an LLM. We first prompted the LLM to revise the reasoning steps (in particular, the formulas) in the explanation text into a normalized form. Then, we prompted the LLM to correct minor errors within the normalized explanations, including formatting issues in formula annotations and inaccuracies in the parameters used during calculation. In the following, we elaborate this two-stage process. More details, including the prompts, are provided in Appendix A.1.1.
@@
We introduced each question with its original explanation and answer to guide the LLM through a few-shot prompt to normalize the explanation. We observed that the quality of the normalized explanations was generally satisfactory. We also required the LLM to present the symbol, numerical value, and unit of each parameter in formulas in the form of a table, as illustrated in Figure 1.
@@
We checked the correctness of the formula format in the explanations by rules, including whether there were omissions in parameter names, symbols, or units, and these minor issues were correctable. To assess the correctness of each normalized explanation, we extracted formulas from the explanation and calculated an answer using the Numbat calculator [^2] https://numbat.dev. Numbat is designed for scientific computation with support for physical units.. If we detected that the LLM generated an ill-formatted or incorrect normalized explanation, we used a few-shot prompt to identify and correct these specific errors. We observed that the questions that remained incorrect despite multiple attempts were of inherently poor quality, e.g., missing important reasoning steps, as illustrated in Appendix A.1.2. These questions were removed from our dataset. After that, our dataset contains 5,420 questions remaining.
@@ 3.3 Formula Database Construction
We divided the construction process of a consolidated formula database into three steps: 1) Merge formulas through symbolic rules. 2) Merge formulas through a semantics-based method. 3) Manually review and correct errors. In Table 3, we present the initial number of formulas and the remaining number of formulas after each step.
@@
This was achieved by comparing formula structures and symbols. Consider the following as an example of judging whether two formulas have the same structure. The three formulas $f_{1}:~a_{1}=(b_{1}+c_{1})/d_{1}$”, $f_{2}:a_{2}=(b_{2}+c_{2})/d_{2}$”, and $f_{3}:b_{1}=a_{1}*d_{1}-c_{1}$” have the same structure because $f_{2}$ can be derived from $f_{1}$ by renaming parameters, and $f_{3}$ can be obtained from $f_{1}$ by transformation. Moreover, in physics, physical quantities are conventionally represented by specific symbols. For example, the mass of an object is often denoted by “mm” and the density of an object is frequently represented by $\rho$”. Subscripts are then used to specify which specific object a physical quantity refers to, such as $\rho_{water}$” for the density of water. Therefore, for each pair of formulas, we first computed all possible transformations of each formula to obtain a set of all its variants. Then, we compared the formula structures in the two sets to determine if the two formulas shared a structure. If so, we checked whether their symbols, with subscripts removed, were identical. If so, we considered these two formulas to be mergeable. When merging, we retained the parameter with the shorter name of the two. After merging based on such symbolic rules, we reduced the number of formulas in the formula database from 12,906 to 1,163.
@@ 3.4 Question Alteration
[…]
@@
To verify that the modified questions were solvable, we used SymPy [^4] https://www.sympy.org/ to solve the equations. During the verification process, some equations could not be solved (e.g., SymPy cannot handle “%” and “1000kg/t”), so we removed these questions. The final number of the remaining questions is 5,324.
@@ 3.5 English Version of FormulaReasoning
We employed the LLM to translate Chinese questions, explanations, and formula database into English. The translation prompts are provided in Appendix A.1.4. To evaluate translation quality, we manually translated a sample of 50 questions and calculated the perplexity of the LLM-translated version and the manually translated version using Qwen models for comparison. We computed the perplexity of LLM and human translations using the Qwen2.5-14B-Instruct model, resulting in scores of $6.28$ and $7.26$, respectively. These findings demonstrate that the current English version maintains a reasonable level of fluency.
We also manually compared the Chinese and translated English versions of the 50 sampled questions. We found that no question had translation errors that would affect question answering.